Find the ratio of: 
 15 kg to 210 g
Find the ratio of: 
 15 kg to 210 g

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

1 kg = 1000 g
Then,
15 kg = 15 × 1000 = 15000 g
Now we have to find the ratio,
= 15000/210
= 1500/21 ... [∵divide both by 3]
= 500/7
So, the required ratio is 500: 7

Thanks